OpenStudy (anonymous):

prova that root5 plus root7 is a irrational number

OpenStudy (ikram002p):

assume its rational then \(\large \sqrt 5+ \sqrt 7=\frac{a}{b} \) \(\large b\sqrt 5+ b\sqrt 7=a \) hence both \(b\large \sqrt 5 \) is rational and \(b\large \sqrt 7 \) is rational

OpenStudy (rational):

using rational root test : \(x^2 - 5\) has no rational roots \(\implies \sqrt{5}\) is irrational \(x^2 - 7\) has no rational roots \(\implies \sqrt{7}\) is irrational since sum of two irrational numbers is irrational, \(\sqrt{5} + \sqrt{7}\) is irrational.

OpenStudy (ikram002p):

two irrational does not apply the sumation irrational like 1/pi - 1/pi =0

OpenStudy (rational):

*sum of two positive irrational numbers

OpenStudy (ikram002p):

the contradiction u got , was since sqrt 5 irrational and b is integer and we assumed \(b \sqrt 5 \) is integer , but we got \(b \sqrt 5 \) ( integer * rational \(\neq \) integer ) which is a contradiction